# ORBIT: Oak Ridge Base Foundation Model for Earth System Predictability

### v3

@@ I Problem Overview
We develop ORBIT, Oak Ridge Base Foundation Model for Earth System Predictability. We designed ORBIT to (1) enable scaling the model size up to 113 billion parameters with throughput up to 1.6 exaFLOPS, (2) incorporate 91 channels of climate variables, and (3) pre-train on 10 different CMIP6 datasets [6] with 1.2 million observation data points. In comparison, the current state-of-the-art ClimaX foundation model uses 115 million parameters, 48 variables and 5 CMIP6 datasets. Our reported 113-billion model represents the largest dense ViTs to date, surpassing the existing largest dense vision model [8] by five-fold, and a thousand-fold larger than the largest AI foundation model for climate—ClimaX [7] and two hundred times larger than domain-specific Stormer model [4]. ORBIT achieves 1.6 exaFLOPS/684 PFLOPS for the 10/113 billion parameter models on 49,152 Frontier AMD GPUs using mixed-single BFLOAT16 precision.
@@ V-F Fine-Tuning Model Prediction
Figure 9: Latitude-Weighted Anomaly Correlation Coefficient (wACC) scores for ORBIT (115 million parameters), ClimaX (same model size), Stormer, FourCastNet, and IFS models in forecasting four atmospheric variables at 1, 14, and 30-day intervals for the testing year 2020.
Fig. 9 illustrates the performance of the 115 million-parameter ORBIT model in forecasting four key atmospheric variables at 1, 14, and 30-day intervals using the ERA5 fine-tuning dataset. This performance is compared with ClimaX [7], Stormer [4], FourCastNet [1], and IFS [28] using the wACC accuracy score. ClimaX is an AI foundation model providing forecasts up to 30 days, and its wACC performance scores were obtained from Fig. 6 of the ClimaX paper [7]. In contrast, Stormer and FourCastNet are not foundation models, but are task-specific AI models trained on ERA5 data with 1.40625 degree and 0.25 degree spatial resolutions, respectively. Their wACC performance scores were obtained from Fig. 9 of the Stormer paper [4] and Fig. 6 of the FourCastNet paper [1]. IFS is a numerical model from the European Center for Medium-Range Weather Forecasting, with its wACC performance scores obtained from Fig. 6 of the ClimaX paper [7]. Notably, Stormer provides forecasts up to 14 days, while FourCastNet and IFS do not offer either 14 or 30-day forecasts.
Fig. 9 highlights the strength of ORBIT, especially evident at longer lead times, where it consistently excels and outperforms ClimaX, Stormer, and IFS. For a 14-day forecast, ORBIT shows improvements of up to 52% over IFS and 166% over Stormer. For the 30-day forecast, ORBIT performs up to 9% better than ClimaX. This superior performance at extended lead times highlights ORBIT’s robustness and reliability in long-term forecasting. For 1-day forecasts, ORBIT’s accuracy is comparable to that of IFS, Stormer, and FourCastNet, and Climax on almost all variables except u10. ORBIT shows a 0.1% better wACC on t850 prediction compared to Climax, Stormer, and FourCastNet. It also shows a 0.27%-1.30% improvement in t2m predictions compared to all others, except IFS. For u10 forecast, ORBIT achieves a 1.26% better wACC than Climax.
It is important to note that both FourCastNet and Stormer were trained and fine-tuned on ERA5 data, with FourCastNet using a higher spatial resolution of 0.25 degree. ClimaX was fine-tuned on each atmospheric variable separately with specifically tailored hyperparameters. In contrast, ORBIT was pre-trained on CMIP6 model simulation data and fine-tuned on ERA5 data at a lower resolution of 1.40625 degree, predicting all four atmospheric variables together as a single task. Despite these differences, ORBIT still demonstrates competitive performance in 1-day forecasts and superior performance in longer-time forecasting compared to other state-of-the-art models.
@@ VI Implications
The ORBIT model’s effective scaling up to 113 billion parameters on a non-NVIDIA platform, such as the Frontier supercomputer, and with up to 1.6 exaFLOP sustained computing throughput also marks a considerable advance in the use of HPC resources for complex AI tasks. With up to 113 billion parameters, ORBIT represents the largest dense vision transformer to date, surpassing the existing largest dense vision transformer by five-fold [8], and a thousandfold larger than the largest foundational climate model, ClimaX [7], and two hundred times larger than domain-specific Stormer model [4]. This success demonstrates that the innovation in scaling algorithms can overcome challenges such as less advanced software stacks and limited interconnect capabilities. It also promotes a more inclusive approach to HPC system development that does not depend exclusively on one type of hardware/software stack, thereby enabling diversity in hardware utilization and encouraging the development of more adaptable and robust HPC systems.
@@ Acknowledgments
The authors thank Verónica G. Melesse Vergara, Mallikarjun (Arjun) Shankar and Bronson Messer for their support of high performance computing resources. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory (ORNL), which is supported by the Office of Science of the U.S. Department of Energy (DOE). This research was primary supported by the AI Initiative Project sponsored by the Director’s Research and Development Program at ORNL, additionally supported by the BER-ASCR SciDAC Program in the DOE, and by Dan Lu’s DOE Early Career Project sponsored by the DOE BER program. This manuscript has been authored by ORNL, operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S.Department of Energy. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.

### v4

@@ I Problem Overview
We develop ORBIT, Oak Ridge Base Foundation Model for Earth System Predictability. We designed ORBIT to (1) enable scaling the model size up to 113 billion parameters with throughput up to 1.6 exaFLOPS, (2) incorporate 91 channels of climate variables, and (3) pre-train on 10 different CMIP6 datasets [6] with 1.2 million observation data points. In comparison, the current state-of-the-art ClimaX foundation model uses 115 million parameters, 48 variables and 5 CMIP6 datasets. Our reported 113-billion model represents the largest dense ViTs to date, surpassing the existing largest dense vision model [8] by five-fold, and a thousand-fold larger than the largest AI foundation model for climate—ClimaX [7] and two hundred times larger than task-specific Stormer model [4]. ORBIT achieves 1.6 exaFLOPS/684 PFLOPS for the 10/113 billion parameter models on 49,152 Frontier AMD GPUs using mixed-single BFLOAT16 precision.
@@ V-F Fine-Tuning Model Prediction
Figure 9: Latitude-Weighted Anomaly Correlation Coefficient (wACC) scores for ORBIT (115 million parameters), ClimaX (same model size), FourCastNet, and IFS models in forecasting four atmospheric variables at 1, 14, and 30-day intervals for the testing year 2020.
Fig. 9 illustrates the performance of the 115 million-parameter ORBIT model in forecasting four key atmospheric variables at 1, 14, and 30-day intervals using the ERA5 fine-tuning dataset. This performance is compared with ClimaX [7], FourCastNet [1], and IFS [28] using the wACC accuracy score. ClimaX is an AI foundation model providing forecasts up to 30 days, and its wACC performance scores were obtained from Fig. 6 of the ClimaX paper [7]. In contrast, FourCastNet is not foundation model, but is task-specific AI models trained on ERA5 data with 0.25 degree spatial resolution. The wACC performance scores were obtained from Fig. 6 of the FourCastNet paper [1]. We did not compare with Stormer [4] since Stormer used different climatology data and a direct comparison is not available. IFS is a numerical model from the European Center for Medium-Range Weather Forecasting, with its wACC performance scores obtained from Fig. 6 of the ClimaX paper [7]. Notably, both ClimaX and ORBIT provides forecasts up to 30 days, while FourCastNet does not offer either 14 or 30-day forecasts.
Fig. 9 highlights the strength of ORBIT, especially evident at longer lead times, where it consistently excels and outperforms ClimaX and IFS. For a 14-day forecast, ORBIT shows improvements of up to 52% over IFS. For the 30-day forecast, ORBIT performs up to 9% better than ClimaX. This superior performance at extended lead times highlights ORBIT’s robustness and reliability in long-term forecasting. For 1-day forecasts, ORBIT’s accuracy is comparable to that of IFS, FourCastNet, and ClimaX on almost all variables except u10. ORBIT shows a 0.1% better wACC on t850 prediction compared to ClimaX and FourCastNet. It also shows a 0.27%-1.30% improvement in t2m predictions compared to all others, except IFS. For u10 forecast, ORBIT achieves a 1.26% better wACC than ClimaX.
It is important to note that the FourCastNet was trained and fine-tuned on ERA5 data and used a higher spatial resolution of 0.25 degree. ClimaX was fine-tuned on each atmospheric variable separately with specifically tailored hyperparameters. In contrast, ORBIT was pre-trained on CMIP6 model simulation data and fine-tuned on ERA5 data at a lower resolution of 1.40625 degree, predicting all four atmospheric variables together as a single task. Despite these differences, ORBIT still demonstrates competitive performance in 1-day forecasts and superior performance in longer-time forecasting compared to other state-of-the-art models.
@@ VI Implications
The ORBIT model’s effective scaling up to 113 billion parameters on a non-NVIDIA platform, such as the Frontier supercomputer, and with up to 1.6 exaFLOP sustained computing throughput also marks a considerable advance in the use of HPC resources for complex AI tasks. With up to 113 billion parameters, ORBIT represents the largest dense vision transformer to date, surpassing the existing largest dense vision transformer by five-fold [8], and a thousandfold larger than the largest foundational climate model, ClimaX [7], and two hundred times larger than task-specific Stormer model [4]. This success demonstrates that the innovation in scaling algorithms can overcome challenges such as less advanced software stacks and limited interconnect capabilities. It also promotes a more inclusive approach to HPC system development that does not depend exclusively on one type of hardware/software stack, thereby enabling diversity in hardware utilization and encouraging the development of more adaptable and robust HPC systems.
@@ Acknowledgments
The authors thank Verónica G. Melesse Vergara, Mallikarjun (Arjun) Shankar and Bronson Messer for their support of high performance computing resources. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory (ORNL), which is supported by the Office of Science of the U.S. Department of Energy (DOE). This research was primary supported by the ORNL’s AI Initiative sponsored by the Director’s Research and Development Program at ORNL, additionally supported by the BER-ASCR SciDAC Program in the DOE, and by Dan Lu’s DOE Early Career Project sponsored by the DOE BER program. This manuscript has been authored by ORNL, operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S.Department of Energy. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.